# Bayesian Density-Density Regression with Application to Cell-Cell Communications

### v1

@@ 3.3 Simulation
Figure 2: Mean in-sample relative predictive errors (7) (fitted computed with $\mathcal{S}$ and predicted computed with $\mathcal{S}^{*}$), and posterior mean squared error $\|A-A^{\star}\|_{F}^{2}+\|b-b^{\star}\|_{2}^{2}$ with corresponding 95% credible interval.
@@
where $\epsilon_{0}\sim\mathcal{N}([0,0]^{\top},I_{2})$, and $m=100$. We consider two simulation scenarios. In scenario 1, we sample $x_{i1},\ldots,x_{im}\overset{i.i.d.}{\sim}\mathcal{N}(\mu_{i},\Sigma_{i})$ with $\mu_{i}\sim\mathcal{N}([0,0]^{\prime},4I_{2})$ and $\Sigma_{i}\sim IW(I_{2},6)$ (inverse Wishart distribution with scale matrix $I_{2}$ and degrees of freedom 6) for $i=1,\ldots,N$. Under scenario 2, we sample $x_{i1},\ldots,x_{im}\overset{i.i.d.}{\sim}\frac{1}{3}\mathcal{N}(\mu_{i1},\Sigma_{i1})+\frac{1}{3}\mathcal{N}(\mu_{i2},\Sigma_{i2})+\frac{1}{3}\mathcal{N}(\mu_{i3},\Sigma_{i3})$, with $\mu_{i1}\sim\mathcal{N}([0,0]^{\top},4I_{2})$, $\mu_{i2}\sim\mathcal{N}([8,8]^{\top},4I_{2})$, $\mu_{i3}\sim\mathcal{N}([-8,8]^{\top},4I_{2})$, and $\Sigma_{i1},\Sigma_{i2},\Sigma_{i3}\sim IW(I_{2},6)$ for $i=1,\ldots,N$.
We compare the proposed Bayesian DDR with Bayesian multivariate linear regression (MLR) replacing the distributions ${G}_{i}$ and ${F}_{i}$ by sample-specific average gene expressions as would be observed in a pseudo-bulk gene expression dataset. That is, the data are $\bar{\mathcal{S}}=\{(\bar{x}_{i},\bar{y}_{i})\}$ where $\bar{x}_{i}=\frac{1}{m_{i1}}\sum_{j=1}^{m_{i1}}x_{ij}$ and $\bar{y}_{i}=\frac{1}{m_{i2}}\sum_{j=1}^{m_{i2}}y_{ij}$. Let $\bar{X}=(\bar{x}_{1},\ldots,\bar{x}_{N})$ and $\bar{Y}=(\bar{y}_{1},\ldots,\bar{y}_{N})$. The Bayesian MLR model is
@@
$\Sigma\mid\bar{Y},\bar{X},A\sim IW(V_{N},\nu_{N}),\quad A\mid\bar{Y},\bar{X},\Sigma\sim\mathcal{MN}(B_{N},\Lambda_{N}^{-1},\Sigma)$
@@
For both, Bayesian DDR and Bayesian MLR, we save 1000 MCMC samples. For the MALA algorithm of Bayesian DDR, we use a step size $\eta=10^{-5}$, and we use $w=10$ in the generalized likelihood. We discard the first 500 samples as burn-in samples. With the final $T=500$ posterior samples $\{A_{1},\dots,A_{T};b_{1},\dots,b_{T}\}$ we evaluate distributional predictions $\bar{G}(z)$ for any predictor distribution ${F}=\frac{1}{m}\sum_{i=1}^{m}\delta_{x_{i}}$ as $\bar{G}(z)=\mathbb{E}[(f_{A,b}\sharp{F})(z)\mid\mathcal{S}]$.
@@
For an assessment of inference as a function of sample size, we evaluate the change in RPE and the change in posterior expected squared error for parameters, $||A-A^{\star}||^{2}_{F}+||b-b^{\star}||^{2}_{2}$, as functions of sample size $N$. Results are shown in Figure 2. To compute RPE, we define the reference model $f_{0}$ in (7) as intercept model, with $A=0$, leaving only the intercept $b$. From Figure 2, we see that both, RPE (in blue) and parameters mean squared error decrease with $N$ as expected. From this simulation, we observe that the generalized posterior under Bayesian DDR already concentrates around the true parameters for theoretically feasible sample sizes $N$.
@@ 4 Example: Cell to Cell Communication
Cell-cell communication plays a fundamental role in understanding complex biological systems, as interactions between cells drive various physiological and pathological processes. A natural way to quantify these interactions is by modeling how the distribution of ligands of one cell type influences the distribution of receptors of another cell type. We demonstrate the proposed generalized Bayesian DDR with an application to inference on the communication between B cells and T cells using the population-scale single-cell data, OneK1K (Yazar et al., , 2022). We select 439 donors who have at least 100 observed cells for both cell types. We formalize cell to cell communication as a regression of a distribution of gene expressions of T cell receptors (as response) on gene expression of B cell ligands (as predictor) to represent “B cell to T cell” communication, and vice versa for “T cell to B cell.” Letting ${G}_{i}$ denote the distribution of T cell receptor gene expression for sample $i$ (as empirical distribution), and ${F}_{i}$ the same for B cell ligands, the regression of T cell on B cell gene expression becomes a density-density regression as in (5) for the pairs $({F}_{i},{G}_{i})$. And similarly for T to B. For the regression of T cell gene expression on B cell gene expression (B cells to T cells) the predictor is the distribution ${F}_{i}$ of B cell ligands {CD40, CD86, ICOSL, IL-6, BAFF, APRIL} and the response is the distribution ${G}_{i}$ of T cell receptors {CD3D, CD3E, CD3G, TRBC1, TRBC2, CD28, ICOS, IL-2R, IFNGR, IL-21R, PD-1, CTLA-4, CXCR5, CCR7}. For T cells to B cells, the predictor is the distribution of T cell ligands {CD40L, IFN-$\gamma$ and the response is the distribution of B cell receptors {IGHM, IGHD, IGHG, IGHA, CD40, ICOSL, IL-21R, IL-6R, BAFFR, CXCR5, CCR7}. The orthogonal nature of the ligand and receptor sets makes the regressions of T on B and B on T complementary. For each task, we split the set of donors into a training set with 80% of the subjects and a test set with the remaining 20%. We standardize all ligands and receptors based on the mean and standard deviation from all cells and all donors in the training set.
@@
For the MALA algorithm, we use a step size $\eta=10^{-4}$. In the definition of the generalized log-likelihood, we use a factor $w=100$, and we use $L=1000$ for approximating the SW distance. We obtain 1000 MCMC samples and discard the first 500 samples for both the Bayesian DDR and the Bayesian MLR. Figure 3 shows the fitted densities $\bar{G}$ of 4 randomly selected donors in the training set (first row of panels) and the predicted densities $\bar{G}$ for 4 random donors in the test set (second row of panels) under the proposed Bayesian DDR of T cell gene expression on B cell gene expression. For comparison, we also indicate the fitted and predicted values for $\bar{y}_{i}$ under the Bayesian MLR. For the Bayesian DDR, we show contour plots of fitted and predicted $\bar{G}=\mathbb{E}[(f_{A,b}\sharp{F})\mid{F},\mathcal{S}]$ for predictor ${F}_{i}={F}$.
[…]
@@ 5 Application to Graph Discovery
To define a weighted fully connected graphs we start by setting up a Bayesian DDR model for each edge $e\in\mathcal{E}_{0}$. Using inference under the DDR model, we evaluate the relative predictive error as defined in (6),
@@
using posterior samples $f_{e1},\ldots,f_{eT}\sim p(f_{e}\mid\mathcal{S}_{e})$ and $f_{e01},\ldots,f_{e0T}\sim p(f_{e0}\mid\mathcal{S}_{e})$ under a reference model $f_{e0}$. For example, if $f_{e}$ is a linear mapping, as we shall use it later, the reference model $f_{e0}$ is the intercept model. Given the posterior mean RPE for all pairs, we report a fully connected graph with the estimated RPEs as weights for directed edges.
For an alternative approach, we define a sparse graph based on FDR and FNR control. We augment the inference model to a decision problem by adding a loss function. First, we use a linear mapping $f_{A,b}(x)=Ax+b$ with parameters $\phi=(A,b)$. We place the horseshoe prior on $A$, as before, in (12). Let $A_{e}$ be the linear coefficient matrix in $f_{A,b}$ that maps from ${F}_{ei}$ to ${G}_{ei}$ for $i=1,\ldots,N_{e}$. We define the $\epsilon$-inclusion probability ($\epsilon>0$):
$p_{e}=\mathbb{P}\Big{(}\max_{ij}\{|A_{eij}|\}>\epsilon~{}\Big{|}~{}\mathcal{S}_{e}\Big{)}\quad i=1,\ldots,d_{e1};\;j=1,\ldots,d_{e2},$ (16)
where $d_{e1}$ and $d_{e2}$ are the dimensions of $\hat{G}_{ei}$ and $\hat{F}_{ei}$. The probability can be estimated using the MCMC samples for $A_{e}$. We interpret the event $\max_{ij}\{|A_{eij}|\}>\epsilon$ as the presence of edge $e$, and use $p_{e}$ to define a parametrized decision rule. We include edge $e=(\ell,r)$ if $p_{e}>0.5$. that is, we use the rule $\widetilde{d}_{e}=I(p_{e}>0.5)$. The rule $\widetilde{d}_{e}$ reduces the decision to the choice of a single decision parameter $\epsilon$. We use a loss function based on posterior expected false discovery rate ($\overline{\text{FDR}}$) and false negative rate ($\overline{\text{FNR}}$) to choose $\epsilon$. In short, for a fixed bound on $\overline{\text{FDR}}$, we choose $\epsilon$ to minimize $\overline{\text{FNR}}$. The two error rates are defined as
@@
where $|\mathcal{E}_{0}|$ is the total number of possible edges (and $0.001$ in the denominator avoids 00-denominators). See, for example, Müller et al., (2004) for a discussion of $\overline{\text{FDR}}$ and $\overline{\text{FNR}}$. We choose $\epsilon$ to minimize $\overline{\text{FNR}}$ subject to $\overline{\text{FDR}}\leq 0.10$.
@@ 6.1 Results for the OneK1K single-cell data
NK Cells to monocytes 102 IL-15, IFN-$\gamma$, FasL IL-12R, IL-15R, IFN-$\gamma$R, CXCR3, Fas, CD40, PD-1, Tim-3
NK Cells to T Cells 700 IL-15, IFN-$\gamma$, CD40L, FasL IL-12R, IL-15R, IFN-$\gamma$R, CD28, ICOS, PD-1, Tim-3
T Cells to NK Cells 700 IFN-$\gamma$, CD86 IL-2R, IL-4R, IL-21R, IFN-$\gamma$R, CCR1, CXCR3, CD28, ICOS, PD-1, Tim-3, NKG2D
@@
NK Cells to B Cells 319 IL-15, IFN-$\gamma$, CXCL8, CD40L IL-2R, IL-15R, IFN-$\gamma$R, CCR1, CXCR2, CD40, ICOS, PD-1, Tim-3
@@
We consider four cell types from the OneK1K dataset: B cells, T cells, NK cells, and monocytes, and aim to infer a graph to represent cell-cell communication between these four subtypes. For each ordered pair of cell types, we extract the ligand genes from the “sender” cell and the receptor genes from the “receiver” cell and regress the distribution of gene expression for the latter onto the distribution of the former. The data are the observed distributions of single-cell RNA-seq gene expression for donors $i=1,\ldots,N$. We select only donors who have at least 100 cells. Table 1 shows the number of donors $N$ for each ordered pair of cell types and the list of ligands and receptors. We standardize all ligands and receptors across all cells and donors. We implement Bayesian DDR using a linear function $f_{A,b}(x)=Ax+b$ and Bayesian MLR. Under both models, we generate 1000 MCMC samples with the first 500 samples discarded as burn-in. We use a step size of $\eta=10^{-4}$ for the MALA algorithm, and a scale factor of $w=100$ in the generalized log-likelihood function.
@@
As before, for visualization we use PCA on the response distribution of each donor to obtain the projection matrix in two dimensions. Figures 7-10 in Supplementary Materials B shows fitted densities $\bar{G}$ under Bayesian DDR for ordered pairs of cell types. For comparison, as before, we also include fitted means under Bayesian MLR. Bayesian DDR reports entire fitted densities $\bar{G}$. In contrast, Bayesian MLR provides a single fitted value for $\bar{y}$, stopping short of representing the entire distributional information of the response.
Figure 4a shows an estimated graph of cell-cell communication, with edge weights based on mean RPE. Lower mean RPE means a higher level of association. We use an exponential kernel based on mean RPE to control the thickness of the edges. We refer the reader to Figure 8 in Supplemental Materials B for boxplots of RPE for all ordered pairs of cell types. For an alternative graph estimate, we show in Figure 4b and Figure 4c the graph based on FDR and FNR control as discussed before. For comparison, we also report an estimated graph based on inference under Bayesian MLR (using the same FDR and FNR control). For Bayesian MLR, we found the optimal threshold $\epsilon=0.858$, yielding posterior expected FDR and FNR of $(\overline{\text{FDR}},\overline{\text{FNR}})=(0.1,0.039)$. Under the Bayesian DDR, the optimal threshold is $\epsilon=0.974$, corresponding to $\overline{\text{FDR}}=0.1,\overline{\text{FNR}}=0.108$. The graph based on Bayesian MLR includes only one edge from B to T cells. Under the Bayesian DDR model we find 6 edges: B to T cells, B to NK cells, monocytes to T cells, monocytes to NK cells, NK cells to monocytes, and monocytes to B cells. Comparing with the mean RPE weights in Figure 4a, we note that the estimated graph under FDR and FNR control seems to select edges with the mean RPE smaller than $0.5$.
Besides facilitating the graph construction, another benefit of choosing the linear function $f_{A,b}$ is the opportunity to explore the interaction between ligands and receptors through inference on the corresponding coefficients in $A$. Figures 11-22 in Supplementary Materials B show boxplots for the linear regression coefficients $A$ for all ordered pairs of cell types. From these summaries, one could determine which ligands are significant in predicting specific receptors. We discuss more details in the Supplementary Materials B.

### v2

@@ 3.4 Simulation
Figure 2: Mean in-sample relative predictive errors (9) (fitted computed with $\mathcal{S}$ and predicted computed with $\mathcal{S}^{*}$), and posterior mean squared error $\|A-A^{\star}\|_{F}^{2}+\|b-b^{\star}\|_{2}^{2}$ with corresponding 95% highest credible interval.
@@
where $\epsilon_{0}\sim\mathcal{N}([0,0]^{\top},I_{2})$, and $m=100$. See below for the choice of $x_{ij}$. We aim to examine whether the proposed model is able to recover the true parameters under a well-specified scenario. For easier visualization we consider the bivariate setting. We consider two simulation scenarios. In scenario 1, we sample $x_{i1},\ldots,x_{im}\overset{i.i.d.}{\sim}\mathcal{N}(\mu_{i},\Sigma_{i})$ with $\mu_{i}\sim\mathcal{N}([0,0]^{\prime},4I_{2})$ and $\Sigma_{i}\sim IW(I_{2},6)$ (inverse Wishart distribution with scale matrix $I_{2}$ and degrees of freedom 6) for $i=1,\ldots,N$. Under scenario 2, we sample $x_{i1},\ldots,x_{im}\overset{i.i.d.}{\sim}\frac{1}{3}\mathcal{N}(\mu_{i1},\Sigma_{i1})+\frac{1}{3}\mathcal{N}(\mu_{i2},\Sigma_{i2})+\frac{1}{3}\mathcal{N}(\mu_{i3},\Sigma_{i3})$, with $\mu_{i1}\sim\mathcal{N}([0,0]^{\top},4I_{2})$, $\mu_{i2}\sim\mathcal{N}([8,8]^{\top},4I_{2})$, $\mu_{i3}\sim\mathcal{N}([-8,8]^{\top},4I_{2})$, and $\Sigma_{i1},\Sigma_{i2},\Sigma_{i3}\sim IW(I_{2},6)$ for $i=1,\ldots,N$.
We compare the proposed Bayesian DDR with Bayesian multivariate linear regression (MLR) replacing the distributions ${G}_{i}$ and ${F}_{i}$ by sample-specific average gene expressions as one would observe in a pseudo-bulk gene expression dataset. That is, the data are $\bar{\mathcal{S}}=\{(\bar{x}_{i},\bar{y}_{i})\}$ where $\bar{x}_{i}=\frac{1}{m_{i1}}\sum_{j=1}^{m_{i1}}x_{ij}$ and $\bar{y}_{i}=\frac{1}{m_{i2}}\sum_{j=1}^{m_{i2}}y_{ij}$. Let $\bar{X}=(\bar{x}_{1},\ldots,\bar{x}_{N})$ and $\bar{Y}=(\bar{y}_{1},\ldots,\bar{y}_{N})$. The Bayesian MLR model is
@@
$\Sigma\mid\bar{Y},\bar{X},A\sim IW(V_{N},\nu_{N}),\quad A\mid\bar{Y},\bar{X},\Sigma\sim\mathcal{MN}(B_{N},\Lambda_{N}^{-1},\Sigma)$ (18)
@@
For both, Bayesian DDR and Bayesian MLR, we save 1000 MCMC samples. In the MALA algorithm for posterior simulation under the Bayesian DDR, we use a step size $\eta=10^{-5}$, and we use $w=10$ in the generalized likelihood. We discard the first 500 samples as burn-in samples. With the final $T=500$ posterior samples $\{A_{1},\dots,A_{T};b_{1},\dots,b_{T}\}$ we evaluate distributional predictions $\bar{G}(z)$ for any predictor distribution ${F}=\frac{1}{m}\sum_{i=1}^{m}\delta_{x_{i}}$ as $\bar{G}(z)=\mathbb{E}[(f_{A,b}\sharp{F})(z)\mid\mathcal{S}]$.
@@
[…]
@@ 4 Example: Cell to Cell Communication
Figure 3: Cell-cell communication: the sender cell produces and sends ligands (signaling molecules) that bind to receptors expressed by the receiver cell.
Cell-cell communication plays a fundamental role in understanding complex biological systems, as interactions between cells drive various physiological and pathological processes. See Figure 3 for an illustration of a sender cell sending signals (ligands), which are received by a receiver cell via the expressed receptors. A natural way to quantify these interactions is by modeling how the distribution of ligands of one cell type influences the distribution of receptors of another cell type. We demonstrate the proposed generalized Bayesian DDR with an application to inference on the communication between B cells and T cells using the population-scale single-cell data, OneK1K (Yazar et al.,, 2022). We select 439 donors who have at least 100 observed cells for both cell types. We formalize cell to cell communication as a regression of a distribution of gene expressions of T cell receptors (as response) on gene expression of B cell ligands (as predictor) to represent “B cell to T cell” communication, and vice versa for “T cell to B cell.” Letting ${G}_{i}$ denote the distribution of T cell receptor gene expression for sample $i$ (as empirical distribution), and ${F}_{i}$ the same for B cell ligands, the regression of T cell on B cell gene expression becomes a density-density regression as in (6) for the pairs $({F}_{i},{G}_{i})$. And similarly for T to B. For the regression of T cell gene expression on B cell gene expression (B cells to T cells) the predictor is the distribution ${F}_{i}$ of B cell ligands {CD40, CD86, ICOSL, IL-6, BAFF, APRIL} and the response is the distribution ${G}_{i}$ of T cell receptors {CD3D, CD3E, CD3G, TRBC1, TRBC2, CD28, ICOS, IL-2R, IFNGR, IL-21R, PD-1, CTLA-4, CXCR5, CCR7}. For T cells to B cells, the predictor is the distribution of T cell ligands {CD40L, IFN-$\gamma$ } and the response is the distribution of B cell receptors {IGHM, IGHD, IGHG, IGHA, CD40, ICOSL, IL-21R, IL-6R, BAFFR, CXCR5, CCR7}. The orthogonal nature of the ligand and receptor sets makes the regressions of T on B and B on T complementary. For each task, we split the set of donors into a training set with 80% of the subjects and a test set with the remaining 20%. We standardize all ligands and receptors based on the mean and standard deviation from all cells and all donors in the training set.
@@
We implement inference under the proposed Bayesian DDR and MLR models. For the MALA algorithm, we use a step size $\eta=10^{-4}$. In the definition of the generalized log-likelihood, we use a factor $w=100$, and we use $L=1000$ for approximating the SW distance. We obtain 1000 MCMC samples and discard the first 500 samples for both the Bayesian DDR and the Bayesian MLR. Figure 4 shows the fitted densities $\bar{G}$ for 4 randomly selected donors in the training set (first row of panels) and the predicted densities $\bar{G}$ for 4 random donors in the test set (second row of panels) under the proposed Bayesian DDR of T cell gene expression on B cell gene expression. For comparison, we also indicate the fitted and predicted values for $\bar{Y}_{i}$ under the Bayesian MLR (18). For the Bayesian DDR, we show contour plots of fitted and predicted $\bar{G}=\mathbb{E}[(f_{A,b}\sharp{F})\mid{F},\mathcal{S}]$ for predictor ${F}_{i}={F}$.
For each donor, we plot the first two projections under PCA of the atoms of the observed response distribution ${G}_{i}$. As a density-density regression, Bayesian DDR reports an entire distribution, while Bayesian MLR naturally only reports a single point estimate of $\bar{Y}_{i}$. We use RPEs as defined in (9) to assess the fit and prediction. For the regression of T cell gene expression on B cell gene expression shown in Figure 4, under the Bayesian DDR we find posterior mean RPEs of $0.233$ with a 95% credible interval of $[0.232,0.235]$ for the fitted densities in the training set, and $0.217~(0.216\text{ to }0.219)$ for the predicted densities for the test set. For the regression of B cell gene expression on T cell gene expression (not shown), we find the posterior mean RPE of $0.617~(0.616\text{ to }0.619)$ for the training set and $0.606~(0.603\text{ to }0.609)$ for the test set.
@@ 5 Application to Graph Discovery
To define a weighted fully connected graphs we start by setting up a Bayesian DDR model for each edge $e\in\mathcal{E}_{0}$. Using inference under the DDR model, we evaluate the relative predictive error as defined in (8) with approximation of SW,
@@
using posterior samples $f_{e1},\ldots,f_{eT}\sim p(f_{e}\mid\mathcal{S}_{e})$ and $f_{e01},\ldots,f_{e0T}\sim p(f_{e0}\mid\mathcal{S}_{e})$ under a reference model $f_{e0}$. For example, if $f_{e}$ is a linear mapping, as we shall use it later, the reference model $f_{e0}$ could be the intercept-only model. Given the posterior mean RPE for all pairs, we report a fully connected graph with the estimated RPEs as weights for directed edges.
For an alternative approach, we define a sparse graph based on FDR and FNR control. We augment the inference model to a decision problem by adding a loss function. As before, we use a linear mapping $f_{A,b}(x)=Ax+b$ with parameters $\phi=(A,b)$. Again, we place a horseshoe prior on $A$, as in (16) before. Let $A_{e}$ be the linear coefficient matrix in $f_{A,b}$ that maps from ${F}_{ei}$ to ${G}_{ei}$ for $i=1,\ldots,N_{e}$. We define the $\epsilon$-inclusion probability ($\epsilon>0$):
$p_{e}=\mathbb{P}\Big(\overbrace{\max_{ij}\{|A_{eij}|\}>\epsilon}^{R_{e}}~\Big|~\mathcal{S}_{e}\Big)\quad i=1,\ldots,d_{e1};\;j=1,\ldots,d_{e2},$ (21)
where $d_{e1}$ and $d_{e2}$ are the dimensions of $\hat{G}_{ei}$ and $\hat{F}_{ei}$. The posterior probability (21) can be estimated using the MCMC samples for $A_{e}$. We interpret the event $R_{e}$ as true presence of edge $e$, and use the posterior probability $p_{e}$ to define a parametrized decision rule. We include edge $e=(\ell,r)$ if $p_{e}>0.5$. that is, we use the rule $\widetilde{d}_{e}=I(p_{e}>0.5)$. The rule $\widetilde{d}_{e}$ reduces the decision to the choice of a single decision parameter $\epsilon$. A true positive happens if $d_{e}=1_{R_{e}}=1$. We use a loss function based on posterior expected false discovery rate ($\overline{\text{FDR}}$) and false negative rate ($\overline{\text{FNR}}$) to choose $\epsilon$. In short, for a fixed bound on $\overline{\text{FDR}}$, we choose $\epsilon$ to minimize $\overline{\text{FNR}}$. The two error rates are defined as
@@
where $|\mathcal{E}_{0}|$ is the total number of possible edges (and $0.001$ in the denominator avoids 0-denominators). See, for example, Müller et al., (2004) for a discussion of $\overline{\text{FDR}}$ and $\overline{\text{FNR}}$. We choose $\epsilon$ to minimize $\overline{\text{FNR}}$ subject to $\overline{\text{FDR}}\leq 0.10$.
@@ 6.1 Results for the OneK1K single-cell data
[…]
@@
[…]
@@
We consider four cell types from the OneK1K dataset: B cells, T cells, NK cells, and monocytes, and aim to infer a graph to represent cell-cell communication between these four subtypes. For each ordered pair $e$ of cell types, we extract the ligand genes from the “sender” cell and the receptor genes from the “receiver” cell and regress the distribution $G_{ei}$ of gene expression for the latter onto the distribution $F_{ei}$ of the former. The data are the observed distributions of single-cell RNA-seq gene expression for donors $i=1,\ldots,N$. We select only donors who have at least 100 cells. Table 1 shows the number of donors $N$ for each ordered pair of cell types and the list of ligands and receptors. We standardize all ligands and receptors across all cells and donors. As before, we implement Bayesian DDR using a linear function $f_{A,b}(x)=Ax+b$ and Bayesian MLR. Under both models, we generate 1000 MCMC samples with the first 500 samples discarded as burn-in. We use a step size of $\eta=10^{-4}$ for the MALA algorithm, and a scale factor of $w=100$ in the generalized log-likelihood function.
@@
As before, for visualization we use PCA on the response distribution of each donor to obtain the projection matrix in two dimensions. Figures 8-11 in Supplementary Materials C shows fitted densities $\bar{G}$ under Bayesian DDR for ordered pairs of cell types. For comparison, as before, we also include fitted means under Bayesian MLR. Bayesian DDR reports entire fitted densities $\bar{G}$. In contrast, Bayesian MLR provides a single fitted value for $\bar{y}$, stopping short of representing the entire distributional information of the response.
Figure 5a shows an estimated graph of cell-cell communication, with edge weights based on mean RPE. Lower mean RPE means a higher level of association. We use an exponential kernel based on mean RPE to control the thickness of the edges. We refer the reader to Figure 9 in Supplemental Materials C for boxplots of RPE for all ordered pairs of cell types. For an alternative graph estimate, we show in Figure 5b and Figure 5c the graph based on FDR and FNR control as discussed before. For comparison, we also report an estimated graph based on inference under Bayesian MLR (using the same FDR and FNR control). For Bayesian MLR, we found the optimal threshold $\epsilon=0.858$, yielding posterior expected FDR and FNR of $(\overline{\text{FDR}},\overline{\text{FNR}})=(0.1,0.039)$. Under the Bayesian DDR, the optimal threshold is $\epsilon=0.974$, corresponding to $\overline{\text{FDR}}=0.1,\overline{\text{FNR}}=0.108$. The graph based on Bayesian MLR includes only one edge from B to T cells. Under the Bayesian DDR model we find 6 edges: B to T cells, B to NK cells, monocytes to T cells, monocytes to NK cells, NK cells to monocytes, and monocytes to B cells. Comparing with the mean RPE weights in Figure 5a, we note that the estimated graph under FDR and FNR control seems to select edges with the mean RPE smaller than $0.5$.
Besides facilitating the graph construction, another benefit of choosing the linear function $f_{A,b}$ is the opportunity to explore the interaction between ligands and receptors through inference on the corresponding coefficients in $A$. Figures 12-23 in Supplementary Materials C show boxplots for the linear regression coefficients $A$ for all ordered pairs of cell types. From these summaries, one could determine which ligands are significant in predicting specific receptors. We discuss more details in the Supplementary Materials C.